a ramp in san francisco at the hotel nikko rises 5.6 inches over a horizontal span of 4'7" solve using pythagorean theorem,(step by step).
Answer by rfer(16322)

a^2+b^2=c^2
5.6^2+55^2=c^2
31.36+3025=3056.36
c=55.28 in = 4'7.2"